The demand function of a good is Q = 100 – 2p. What is the elasticity at the point p=10 and Q=800?
Answer #1

Puye No Date hion dp go goo
Answer #2
Elasticity of demand=-p/q×dq/dp. (d/dp of 100-2p=-2) - P/100-2p×-2...we have p is 25 at MR=o so 50/50=1
source: Textbook
answered by: anonymous